OpenStudy (jack1):

hey guys, have a ... lets go with "difficult" question: if x*y = P and x = a* sin (phi) and y = b* sin (phi - theta) and a = 162 and b = 7.5 and theta = 50 degrees ... how do i find the value for P(maximum) ... please? and also the corresponding value for phi i guess?

OpenStudy (jack1):

hey @tkhunny how u dooin? any ideas on the above dude?

OpenStudy (tkhunny):

Have you considered substituting everything and finding all the 1st and 2nd partial derivatimes, \(\dfrac{dP}{d\phi}\)?

OpenStudy (jack1):

i did, but then i realised i learned partial derivatives last term, so over the course of 4 months have forgotten how to do them... dammit

OpenStudy (jack1):

know a good refresher site to relearn partial derivatives dude?

OpenStudy (tkhunny):

We can simply remove the \(\theta\) problem with trig identities. \(\sin(\phi - \theta) = \sin(\phi)\cos(\theta)-\cos(\phi)\sin(\theta)\)

OpenStudy (tkhunny):

Actually, there is no PARTIAL derivative problem, here. \(\theta\) is a constant.

OpenStudy (jack1):

yup: 50

OpenStudy (tkhunny):

Okay, we're down to \(\dfrac{d}{d\phi}[1215\sin(\phi)\sin(\phi - 50º)]\). What do you get for that derivative?

OpenStudy (jack1):

1215 [cos (phi-50) sin(phi)] + cos (phi) sin (phi + 50) ... is that right?

OpenStudy (tkhunny):

You have some trigonometry to go to make your life easier, unless you're okay with that result. With that correction, It's right, but it's way ugly. I get \(\dfrac{dP}{d\phi} = 1215\cdot\sin\left(2\phi - \dfrac{5\pi}{18}\right) = 2430\cdot\sin^{2}\left(\phi + \dfrac{\pi}{9}\right) - 1215\) Kind of remarkable that those are the same.

OpenStudy (jack1):

i dont think im doing it right, the answer's supposedly 1000 Watts @ phi = 115 degrees cheers anyway @tkhunny for the help, i think i needs more study time to understand the basics before i try for the specifics

OpenStudy (jack1):

k i dont understand how they got it, but one of my mates sent me how he did it (its a 2008 practice exam we're going through): if y = -1215. sin (50 - x) sin(x) (x is phi) y' = -1215. sin (50 - 2x)... he didn't email the explanation, just the result so 0 = -1215. sin (50 - 2x) so x = phi = 25 -(pi n /2) = so that'll give me the low and the high points, -65 was a minimum, 115 was a maximum it seems to be right, (P = 998 Watts) but I clearly don't get the maths (derivatives using product rule then chain rule consecuatively ... or spelling consecuatively...) so will have to work on it whay more. Thanks again @tkhunny Jack

OpenStudy (tkhunny):

You don't recognize one of my derivatives? " y' = -1215. sin (50 - 2x)"